What capacitance is needed to store 3.00μC of charge at a voltage of 120 V?



1
Expert's answer
2022-02-10T15:02:23-0500

The capacitance needed to store "3\mu C" of charge when "120" volts is applied can be found by taking the charge divided by the voltage.

That is

"\displaystyle C=\frac{3.00 \mu C\times \left[\left[\frac{10^{-6}C}{1\mu C}\right]\right]}{120V}"


"\displaystyle C=2.50\times10^{-8}F\left\{\frac{1nF}{10^{-9}F}\right\}"


"C=25.0nF"

which is the capacitance needed.